View History Bookmarks People Window Help Login or Regis x owLv2 l online teaching and x G hcr acid name Google Sea x enow.com/ilrn/takeAssignment/takeCovalentActivity do?iocator assignment-take&takeAssignmentsession.... pHouse KylesFileserver LGoogle Drive Oe-campus Sakai URI D owLv2 Other Book This question has multiple parts. Work all the parts to get the most points. Calculate the molar concentration of OH in water solutions with the following H30 molar a H30 0.068 M OH CH30 OH C [H3O 0.0068 M OH Od [Hao 60 x 10 10 M IOH Ge Hao 6.0 x 10-2 M
Answer #1

[H3O+]*[OH-] = 1.0*10-14

[OH-] = 1.0*10-14/[H3O+]

A. [H3O+] = 0.068M

[OH-] = 1.0*10-14/0.068 = 1.47*10-13M

B.[H3O+] = 5.8*10-4M

[OH-] = 1.0*10-14/5.8*10-4 = 1.72*10-11M

C.[H3O+] = 0.0068M

[OH-] = 1.0*10-14/0.0068 = 1.47*10-12M

D.[H3O+] = 6.0*10-10M

[OH-] = 1.0*10-14/6.0*10-10 = 1.67*10-5M

E.[H3O+] = 6.0*10-2M

[OH-] = 1.0*10-14/6.0*10-2 = 1.67*10-13M